# Noisy MRI Reconstruction via MAP Estimation with an Implicit Deep-Denoiser Prior

### v1

@@ I Introduction and Background
Magnetic resonance imaging (MRI) is a widely used, non-invasive technique for the clinical diagnosis of conditions affecting many organs. However, a primary limitation of MRI is its slow acquisition speed, which is constrained by fundamental physical principles. To accelerate scanning, parallel imaging methods use multiple receiver coils to acquire undersampled k-space data, thereby reducing scan time and performing a regularized reconstruction at inference [Rep2Rep2025]. This regularization is traditionally guided by hand-crafted priors, however, recently end-to-end trained deep neural networks have shown state-of-the-art performance [lpds2025].
Another growing body of related literature is zero-shot deep denoising diffusion generative models (DDGMs) [zhu2023denoising, chung2024decomposed, hoDenoisingDiffusionProbabilistic2020, daras2024survey, rozet2024learning, song2023pseudoinverseguided]. These models have emerged as a powerful class of methods for solving image inverse problems. In natural image domains, these models have demonstrated an exceptional ability to produce realistic and high-fidelity reconstructions. Fundamentally, diffusion models are trained to learn the score function, i.e. the gradient of the log-probability of the data distribution, through a denoising process that gradually removes noise from corrupted samples. By reversing this diffusion process, the models can generate samples from the learned data distribution and effectively recover images from incomplete or degraded measurements [daras2024survey]. In MRI, diffusion-based approaches have shown promising results in achieving high acceleration rates under simulated noise conditions [chung2024decomposed]. However, DDGMs are typically formulated using stochastic differential equations or Markov chains, both of which are decoupled from the underlying MRI physics and the applicability of DDGMs to real-world MRI data with measurement noise introduced by the scanner remains largely unexplored.
In a series of recent publications [janjusevicCDLNet2022, janjusevicGDLNet2022, janjusevic2024groupcdl, lpds2025], we have introduced a framework for the interpretable design of deep denoisers that generalize effectively across, and even beyond, their training noise levels. We have further extended this framework to the problem of MRI reconstruction, demonstrating its applicability in both supervised and self-supervised learning settings. In this work, we extend these ideas to diffusion-based models and propose an interpretable construction framework for diffusion models tailored to MRI reconstruction. Specifically, we build upon and expand the formulation of diffusion models introduced in Kadkhodaie et al. [kadkhodaie2021stochastic] to address inverse problems in the presence of observation noise. We show state-of-the-art performance of the proposed approach on both synthetically added noise and real MRI acquisitions with scanner noise.
@@ I-A MRI Reconstruction Problem
We model our observations as a subsampled multicoil Fourier domain signal $\bm{y}\in\mathbb{C}^{N_{s}C}$}, with $C$ coils, of a ground-truth image $\bm{x}\in\mathbb{C}^{N}$}, with fourier-sampling locations denoted by index-set $\Omega$ and coil-sensitivity maps $\bm{s}\in\mathbb{C}^{NC}$},
@@
where $\bm{F}$ is the $N$-pixel 2D-DFT matrix, $\circ$ denotes element-wise multiplication, and $\bm{I}_{\Omega}\in\{0,1\}^{N_{s}\times N}$} is the row-removed identity matrix with kept rows indicated by $\Omega$. For simplicity, we consider our observations to be noise-whitened such that $\bm{\nu}_{c}\sim\mathcal{N}(0,\sigma^{2}\bm{I})$} is per-channel i.i.d. Gaussian noise with noise-level $\sigma$ identical in each channel. We denote this observation model through the encoding operator $\bm{A}$ defined such that $\bm{y}=\bm{A}\bm{x}+\bm{\nu}$}. We say the observation is accelerated by a factor of $R={N/N_{s}}$.
@@ I-B Stochastic Ascent with a Denoiser’s Implicit Prior
In this section, we present the stochastic gradient descent algorithm with a denoiser’s implicit prior of Kadkhodaie et al. [kadkhodaie2021stochastic]. Let $\bm{z}=\bm{x}+\bm{n}$ where $\bm{x}\sim p$ and $\bm{n}\sim g_{\sigma}=\mathcal{N}(0,\sigma^{2}\bm{I})$. Then we may obtain the following relation (Tweedie’s Formula [kadkhodaie2021stochastic]) directly from expressing the noisy image distribution $p_{\sigma}$ as the convolution of the image distribution and the Guassian,
@@
Tweedie’s formula (4) tells us that the MMSE estimate of the ground-truth $\bm{x}$ given a noisy $\bm{z}$ is achieved by a single gradient ascent step on the log-probability of $p_{\sigma}$, $\operatorname{\mathbb{E}}[\bm{x}|\bm{z}]=\bm{z}+\sigma^{2}\nabla_{\bm{z}}\log p_{\sigma}(\bm{z})$} with step-size $\sigma^{2}$. However, $p_{\sigma}$ is only a Gaussian-blurred version of our true image distribution $p$. The key-insight of Kadkhodaie et al. [kadkhodaie2021stochastic] is that we may interpret the score function of the noisy distribution, $\sigma^{2}\nabla_{\bm{z}}\log p_{\sigma}(\bm{z})$}, as an approximate ascent direction on the true image distribution. If we take a fractional step in this direction, our new estimate will be closer to the true image distribution, less noisy ($\sigma\downarrow$), and we will have access to an even more accurate ascent direction w.r.t. the true image distribution. Upon this insight, Kadkhodaie et al. [kadkhodaie2021stochastic] propose the following algorithm, which may be interpreted as a coarse-to-fine gradient ascent on the log-probability,
@@
with starting point $\bm{z}_{0}\sim\mathcal{N}(0,\bm{I})$}, step-size $h_{t}\in[0,1]$} and stochastic noise term $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\gamma_{t}^{2}\bm{I})$} introduced to help escape local maxima. In the same work, Kadkhodaie et al. [kadkhodaie2021stochastic] extended their algorithm to handle conditioning on measurement data, $\bm{y}=\bm{A}\bm{x}$}, by performing maximization of the log-posterior using Baye’s rule, i.e, $\operatorname*{arg\,max\,}_{\bm{x}}\log p(\bm{x}|\bm{y})=\operatorname*{arg\,max\,}_{\bm{x}}\log p(\bm{x})+\log p(\bm{y}|\bm{x})$}. However, their method relied on $\bm{y}$ being noise-free and having an easily SVD decomposable observation operator $\bm{A}$, limiting its direct application to large-scale MR imaging inverse-problems.
@@ II-A The Implicit MAP Algorithm
using the prior implicitly encoded by a trained denoiser $f$ to access the prior score function via Tweedie’s formula (4). We switch to the notation to variable $\bm{z}$ to indicate that our intermediate variables $\bm{z}_{t}$ belong to the noisy image distribution, $\bm{z}_{t}\sim p_{\sigma_{t}}$}. To access the likelihood score-function, $\nabla_{\bm{z}_{t}}\log p(\bm{y}|\bm{z}_{t})$}, we model $\bm{z}_{t}$ as a noisy version of our desired ground-truth signal, $\bm{z}_{t}\sim\mathcal{N}(\bm{x},~\sigma_{t}^{2}\bm{I})$}. Following [song2023pseudoinverseguided, daras2024survey], we approximate the distribution of a denoised $\bm{z}_{t}$, i.e. $f(\bm{z}_{t})$, with a Gaussian distribution (i.e. a Laplace Approximation),
@@
where the variance-matrix $\bm{V}_{t}$ is a modeling parameter which we set to $\bm{V}_{t}=\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}\bm{I}$} in practice, following [song2023pseudoinverseguided]. We derive the likelihood function as,
@@
With $\operatorname{\mathbb{E}}[\bm{x}|\bm{z}_{t}]\approx f(\bm{z}_{t})$, we implement $\nabla_{\bm{z}_{t}}\operatorname{\mathbb{E}}[\bm{x}|\bm{z}_{t}]^{H}\bm{v}\approx\bm{J}^{H}_{t}\bm{v}$ with a vector-jacobian product (VJP) with respect to our trained denoiser ff [^1] The VJP operator is accessible via standard automatic differentiation libraries, often via a pullback function.. This allows us to solve the symmetric linear system in the likelihood term (9) via the conjugate gradient method [song2023pseudoinverseguided]. Lastly, we employ a stochastic noise term ($\bm{\epsilon}_{t}$) with stochasticity term $\beta\in(0,1]$} for purposes of escaping local minima, achieving the stochastic MAP estimation algorithm with implicit denoiser prior (ImMAP), shown in Algorithm 1.
@@ II-B Comparison of ImMAP to Other Diffusion-based Methods
with $\bm{x}_{0}\sim p_{0}$ and $\bm{W}_{t}$ denoting a Wiener process. Solving (10) backwards in time is then interpreted as obtaining a sample from the noise-free image distribution $p_{0}$, for appropriately chosen $h$ and $l$ [daras2024survey]. These equations can be easily extended to sampling from a conditional distribution $p_{0}(\bm{x}|\bm{y})$ given measurements $\bm{y}$ [daras2024survey].
This formulation of DDGMs introduces a plethora of modeling choices/assumptions having solely to do with the SDE/Markov process, unrelated to any imaging inverse problem. In contrast, the derivation of ImMAP makes no such modeling assumptions and instead relies only on Tweedie’s formula (4[daras2024survey] and MRI physics.
Recently, Chung et al. proposed Decomposed Diffusion Sampling (DDS) [chung2024decomposed], a DDGM designed specifically for large scale noisy inverse problems. DDS is formulated as a Krylov-subspace method, relying on a “few-step” conjugate gradient method to keep its intermediate estimates close to a Krylov subspace tangent to the noisy image manifold. However, the existence of such a tangent space is not necessarily valid in practice, and their choice of number of conjugate-gradient steps is shown to greatly influence their algorithm’s performance. Additionally, they introduce another hyperparameter $\gamma$ to control the trade-off between data-consistency and their score-based denoiser, which we find in practice needs tuning for different different MRI acceleration factors.
@@ III-A Dataset and Experimental Settings
The performance of all models was evaluated on an internal MRI brain dataset consisting of T2w multicoil k-space data, 72 fully-sampled datasets were used for training with 8 used as a held-out test set. This data was used in two experimental settings: 1) the simulated measurement setting, where the fully-sampled coil- combined reconstruction serves as ground truth and is subsequently passed through a parallel MR observation operator (allowing us to meaningfully compute quantitative image-quality metrics), and 2) a retrospectively subsampled k-space setting, where the fully-sampled multi-coil k-space is subsampled along a single cartesian readout direction. In the latter case, the noise in each coil makes these measurements deviate from any specific measurement operator, and hence they represent experiments with real Gaussian noise.

### v2

@@ I Introduction and Background
Magnetic resonance imaging (MRI) is a widely used, non-invasive technique for the clinical diagnosis of conditions affecting many organs. However, a primary limitation of MRI is its slow acquisition speed, which is constrained by fundamental physical principles. To accelerate scanning, parallel imaging methods use multiple receiver coils to acquire undersampled k-space data, thereby reducing scan time and performing a regularized reconstruction at inference [4]. This regularization is traditionally guided by hand-crafted priors, however, recently end-to-end trained deep neural networks have shown state-of-the-art performance [6].
Another growing body of related literature is zero-shot deep denoising diffusion generative models (DDGMs) [16, 1, 3, 2, 10, 11]. These models have emerged as a powerful class of methods for solving image inverse problems. In natural image domains, these models have demonstrated an exceptional ability to produce realistic and high-fidelity reconstructions. Fundamentally, diffusion models are trained to learn the score function, i.e. the gradient of the log-probability of the data distribution, through a denoising process that gradually removes noise from corrupted samples. By reversing this diffusion process, the models can generate samples from the learned data distribution and effectively recover images from incomplete or degraded measurements [2]. In MRI, diffusion-based approaches have shown promising results in achieving high acceleration rates under simulated noise conditions [1]. However, DDGMs are typically formulated using stochastic differential equations or Markov chains, both of which are decoupled from the underlying MRI physics and the applicability of DDGMs to real-world MRI data with measurement noise introduced by the scanner remains largely unexplored.
In a series of recent publications [7, 8, 5], we have introduced a framework for the interpretable design of deep denoisers that generalize effectively across, and even beyond, their training noise levels has been introduced. We further extended this framework to the problem of MRI reconstruction, demonstrating its applicability in both supervised and self-supervised learning settings [6]. In this work, we expand these ideas to diffusion-based models and propose an interpretable construction framework for diffusion models tailored to MRI reconstruction. Specifically, we expand upon the formulation of diffusion models introduced in Kadkhodaie et al. [9] to address inverse problems in the presence of observation noise. We show state-of-the-art performance of the proposed approach on both synthetically added noise and real MRI acquisitions with scanner noise.
@@ I-A MRI Reconstruction Problem
We model our observations as an undersampled multicoil Fourier domain signal $\bm{y}\in\mathbb{C}^{N_{s}C}$}, with $C$ coils, of a ground-truth image $\bm{x}\in\mathbb{C}^{N}$}, with Fourier sampling locations denoted by index-set $\Omega$ and coil-sensitivity maps $\bm{s}\in\mathbb{C}^{NC}$},
@@
where $\bm{F}$ is the $N$-pixel 2D-DFT matrix, $\bm{s}_{c}\in\mathbb{C}^{N}~\forall~c$}, $\circ$ denotes element-wise multiplication, and $\bm{I}_{\Omega}\in\{0,1\}^{N_{s}\times N}$} is the row-removed identity matrix with kept rows indicated by $\Omega$. We consider our observations to be noise-whitened such that $\bm{\nu}[n]\sim\mathcal{N}(0,\bm{\Sigma}_{y})$} is Gaussian noise with noise-covariance matrix $\bm{\Sigma}_{y}\in\mathbb{C}^{C\times C}$} identically distributed for all pixels $n\in[1,N_{s}]$}. We denote this observation model through the encoding operator $\bm{A}$ defined such that $\bm{y}=\bm{A}\bm{x}+\bm{\nu}$}. We say the observation is accelerated by a factor of $R={N/N_{s}}$}. For the remainder of this manuscript, we follow the mathematical notation of [4].
@@ I-B Stochastic Ascent with a Denoiser’s Implicit Prior
In this section, we present the stochastic gradient descent algorithm with a denoiser’s implicit prior of Kadkhodaie et al. [9]. Let $\bm{z}=\bm{x}+\bm{n}$ where $\bm{x}\sim p$ and $\bm{n}\sim g_{\sigma}=\mathcal{N}(0,\sigma^{2}\bm{I})$. Then we may obtain the following relation (Tweedie’s Formula [9]) directly from expressing the noisy image distribution $p_{\sigma}$ as the convolution of the noiseless image distribution $p$ and the Gaussian PDF $g_{\sigma}$,
@@
Tweedie’s formula (4) tells us that the MMSE estimate of the ground-truth $\bm{x}$ given a noisy $\bm{z}$ is achieved by a single gradient ascent step on the log-probability of $p_{\sigma}$, $\operatorname{\mathbb{E}}[\bm{x}|\bm{z}]=\bm{z}+\sigma^{2}\nabla_{\bm{z}}\log p_{\sigma}(\bm{z})$} with step-size $\sigma^{2}$. However, $p_{\sigma}$ is only a Gaussian-blurred version of our true image distribution $p$. The key-insight of Kadkhodaie et al. [9] is that we may interpret the score function of the noisy distribution, $\sigma^{2}\nabla_{\bm{z}}\log p_{\sigma}(\bm{z})$}, as an approximate ascent direction on the true image distribution. If we take a fractional step in this direction, our new estimate will be closer to the true image distribution, less noisy ($\sigma\downarrow$), and we will have access to an even more accurate ascent direction w.r.t. the true image distribution. Upon this insight, Kadkhodaie et al. [9] propose the following algorithm, which may be interpreted as a coarse-to-fine gradient ascent on the log-probability,
@@
with starting point $\bm{z}_{0}\sim\mathcal{N}(0,\bm{I})$}, step-size $h_{t}\in[0,1]$} and stochastic noise term $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\gamma_{t}^{2}\bm{I})$} introduced to help escape local maxima. In the same work, Kadkhodaie et al. [9] extended their algorithm to handle conditioning on measurement data, $\bm{y}=\bm{A}\bm{x}$}, by performing maximization of the log-posterior using Baye’s rule, i.e, $\operatorname*{arg\,max\,}_{\bm{x}}\log p(\bm{x}|\bm{y})=\operatorname*{arg\,max\,}_{\bm{x}}\log p(\bm{x})+\log p(\bm{y}|\bm{x})$}. However, their method relied on $\bm{y}$ being noise-free and having an easily SVD decomposable observation operator $\bm{A}$, limiting its direct application to large-scale MR imaging inverse-problems.
@@ II-A The Implicit MAP Algorithm
using the prior implicitly encoded by a trained denoiser $f$ to access the prior score function via Tweedie’s formula (4). We switch to the notation to variable $\bm{z}$ to indicate that our intermediate variables $\bm{z}_{t}$ belong to the noisy image distribution, $\bm{z}_{t}\sim p_{\sigma_{t}}$}. To access the likelihood score-function, $\nabla_{\bm{z}_{t}}\log p(\bm{y}|\bm{z}_{t})$}, we model $\bm{z}_{t}$ as a noisy version of our desired ground-truth signal, $\bm{z}_{t}\sim\mathcal{N}(\bm{x},~\sigma_{t}^{2}\bm{I})$}. Following [11, 2], we approximate the distribution of a denoised $\bm{z}_{t}$, i.e. $f(\bm{z}_{t})$, with a Gaussian distribution (i.e. a Laplace Approximation),
@@
where the variance-matrix $\bm{V}_{t}$ is a modeling parameter which we set to $\bm{V}_{t}=\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}\bm{I}$} in practice, following [11]. We derive the likelihood function as,
@@
With $\operatorname{\mathbb{E}}[\bm{x}|\bm{z}_{t}]\approx f(\bm{z}_{t})$, we implement $\nabla_{\bm{z}_{t}}\operatorname{\mathbb{E}}[\bm{x}|\bm{z}_{t}]^{H}\bm{v}\approx\bm{J}^{H}_{t}\bm{v}$ with a vector-jacobian product (VJP) with respect to our trained denoiser ff [^1] The VJP operator is accessible via standard automatic differentiation libraries, often via a pullback function.. This allows us to solve the symmetric linear system in the likelihood term (9) via the conjugate gradient method [11]. Lastly, we employ a stochastic noise term ($\bm{\epsilon}_{t}$) with stochasticity term $\beta\in(0,1]$} for purposes of escaping local minima, achieving the stochastic MAP estimation algorithm with implicit denoiser prior (ImMAP), shown in Algorithm 1.
@@ II-B Comparison of ImMAP to Other Diffusion-based Methods
with $\bm{x}_{0}\sim p_{0}$ and $\bm{W}_{t}$ denoting a Wiener process. Solving (10) backwards in time is then interpreted as obtaining a sample from the noise-free image distribution $p_{0}$, for appropriately chosen $h$ and $l$ [2]. These equations can be easily extended to sampling from a conditional distribution $p_{0}(\bm{x}|\bm{y})$ given measurements $\bm{y}$ [2].
This formulation of DDGMs introduces a plethora of modeling choices/assumptions having solely to do with the SDE/Markov process, unrelated to any imaging inverse problem. In contrast, the derivation of ImMAP makes no such modeling assumptions and instead relies only on Tweedie’s formula (4[2] and MRI physics.
Recently, Chung et al. proposed Decomposed Diffusion Sampling (DDS) [1], a DDGM designed specifically for large scale noisy inverse problems. DDS is formulated as a Krylov-subspace method, relying on a “few-step” conjugate gradient algorithm to keep its intermediate estimates close to a Krylov-subspace tangent to the noisy image manifold. However, the existence of such a tangent space is not necessarily valid in practice, and their choice of number of conjugate-gradient steps is shown to greatly influence their algorithm’s performance [1]. Additionally, they introduce another hyperparameter $\gamma$ to control the trade-off between data-consistency and their score-based denoiser, which we find in practice needs tuning for different MRI noise-levels (see Section III-D).
@@ III Results
Figure 1: Quantitative performance of end-to-end (LPDSNet), and denoising diffusion based (DDS, ImMAP) reconstruction methods on the T1w brain test dataset. NRMSE/LPIPS/SSIM (%) shown at $8\times$ acceleration and across noise-levels $(\sigma_{y})$. Noise is considered in two scenarios: simulated noise added to the coil-combined fully-sampled image data (i.e. $(\bm{\Sigma}_{y})_{cc}=\sigma_{y}\in(0.01,0.02,0.05)$}), and measurement noise (meas., $\bm{\Sigma}_{y}=\hat{\bm{\Sigma}}_{y}$) where the original fully-sampled k-space data is retrospectively undersampled. Error bars show standard-error of the mean over 16 test-volumes.
Figure 2: Visual comparison of ImMAP joint denoising and reconstruction MRI at $8\times$ acceleration against classical reconstruction (SENSE [13]), end-to-end deep reconstruction (LPDSNet [6]), denoising diffusion model (DDS [1]), across noise-levels $(\sigma_{y})$. Quantitative metrics (NRMSE/LPIPS/SSIM (%)) are shown in yellow above each reconstruction.
Figure 3: Ablation of ImMAP’s noise-injection parameter $\beta\in(0,1]$. Left: average NRMSE over the T1w brain test set vs. $\beta$, across noise-levels $\sigma_{y}=(0.01,0.02,0.05,\text{meas.})$}. Standard error of the mean shown via shaded regions. Right: Corresponding average number of iterations vs. $\beta$, across noise-levels.
Figure 4: DDS performance (average volume NRMSE) on $8\times$ accelerated T1w brain test-set with varying $\gamma$ hyperparameter over different noise levels $\sigma_{y}=(0.01,0.02,0.05,\text{meas.})$}. Standard error of the mean shown via shaded regions. DDS takes 83 iterations.
Figure 5: Visualization of $8\times$ reconstruction at $\sigma_{y}=0.02$ when changing hyperparameters for DDS ($\gamma$) and ImMAP ($\beta$) on a slice from the T1w brain test-set. NRMSE/LPIPS/SSIM $(\%)$ shown on each image as compared to the fully-sampled reconstruction.
@@ III-A Dataset and Experimental Settings
The performance of all models was evaluated on an internal MRI brain dataset consisting of fully-sampled T1w multicoil k-space data (i.e. $\bm{y}=\bm{F}\bm{S}\bm{x}+\bm{\nu}$}). 66 fully-sampled volumes ($\sim$60 slices each) were used for training with 16 used as a held-out test set. This data was used in two experimental settings,
  1. 1.
    simulated noise: the fully-sampled coil-combined image reconstruction ($\bm{x}^{\prime}=\bm{S}^{H}\bm{F}^{H}\bm{y}$}) is passed through an MRI encoding operator $\bm{A}=\bm{I}_{\Omega}\bm{F}\bm{S}$},
    $\bm{y}^{\prime}=\bm{A}\bm{x}^{\prime}+\bm{\nu}^{\prime},$ (11)
    with noise $\bm{\nu}^{\prime}\sim\mathcal{N}(0,\,\sigma_{y}\bm{I})$ added in simulation. Reconstruction involves recovering $\bm{x}^{\prime}$} from k-space $\bm{y}^{\prime}$, which contains measurement coil-noise. Quantitative error measurement compares reconstructions $\hat{\bm{x}}$ to $\bm{x}^{\prime}$ with no irreducible error due to noise. For simulated noise experiments, white Gaussian noise ($\bm{\Sigma}_{y}=\operatorname{\mathbf{diag}}(\sigma_{y})$}) was added under three noise-levels, $\sigma_{y}\in(0.01,0.02,0.05)$}.
  2. 2.
    measurement noise: the fully-sampled multi-coil k-space is retrospectively undersampled,
    $\bm{y}^{\prime}=\bm{I}_{\Omega}\bm{y}=\bm{I}_{\Omega}(\bm{F}\bm{S}\bm{x}+\bm{\nu})=\bm{A}\bm{x}+\bm{I}_{\Omega}\bm{\nu},$ (12)
    with noise $\bm{\nu}$ coming directly from the scanner. Reconstruction involves recovering the ground-truth noise-free image $\bm{x}$ from k-space $\bm{y}^{\prime}$. Note that $\bm{x}$ does not contain measurement noise and is inaccessible to us (due to the SNR of the collected k-space data). Quantitative error measurement compares reconstructions $\hat{\bm{x}}$ to $\bm{x}^{\prime}$} with irreducible error due to subsampled measurement noise $\bm{I}_{\Omega}\bm{\nu}$}, which is inseparable from the measurement k-space data (and hence inseparable from $\bm{x}^{\prime}$). For measurement noise experiments, estimated coil-noise covariance matrices $\hat{\bm{\Sigma}}_{y}$ were obtained via a wavelet-filter multicoil image domain method, following [4]. We denote experiments in the measurement noise setting as $\sigma_{y}=\text{meas.}$” for notational brevity.
For all experiments, $8\times$ Cartesian subsampling was used in a single readout direction with a $4\%$ wide region of k-space included. Coil sensitivity maps were obtained by the ESPIRiT algorithm [13] applied to the center of k-space.